Y=5-2x and Y= 2+3x in the same graph Identify the equilibrium value for x and y

1
Expert's answer
2022-02-15T16:10:32-0500

Put "Y=Y"

"5-2x=2+3x"

"5x=3"

"x=3\/5=0.6"

"Y=5-2x=5-2\times 3/5=5-6/5= 19/5=3.8"